OpenStudy (anonymous):

In a survey of 1000 household,washing machines, vacuum cleaners and refrigerators were counted. Each house has at least one of these appliances, 400 had no refrigerator, 380 had no vacuum cleaners, and 542 had no washing machines. 294 had vacuum cleaner, and washing machine,277 both refrigerator and washing machine. Q.1) How many had only vacuum cleaners? Q.2) How many had only refrigerator? Q.3) How many had at least two of the three appliances? Q.4) How many had all three appliances? Q.5) How many had exactly one appliance?

OpenStudy (anonymous):

Try to reverse all conditions to things people have, if 400 people have no refrigerator, then how many people do? And so on. Try to draw a cannot diagram and fill all the known number on, it might be a bit messy, but it's the clearest way

OpenStudy (anonymous):

I mean vann diagram.

OpenStudy (anonymous):

unable to get the answers through vann diagram...please smbd explain it in briefly

OpenStudy (anonymous):

So if we add the amount people who has refrigerator and people who has vacuum cleaner together, we will find out we counted some people twice. The people who owns both the refrigerator and vacuum cleaner Does that help?

OpenStudy (anonymous):

can you please draw it?

OpenStudy (anonymous):

Sorry, on my phone.

OpenStudy (anonymous):

I heartly wish if you can ansewer it later on..

OpenStudy (anonymous):



OpenStudy (anonymous):

1000-(a+c+d+f)=400 1000-(b+c+d+e)=380 1000-(g+e+f+d)=542 b+c+d+e+g+f=294 a+c+d+f+e+g=277